The coefficient of friction between a 0.17 kg puck and and the ice is μk = 0.15. If the puck leaves a hockey stick traveling at 20 m/s, what is its speed when it reaches the goalie 15 m away? How much time does the goalie have to react? Assume the puck travels on the ice the whole time.
Answer #1

using conservsation of energy

energy lost in friction = change in KE

u mg d = 0.5 m ( v^2 - u^2)

0.15* 9.8* 15 = 0.5* ( 20^2 - u^2)

u = 18.865 m/s

=========

b)

Acceleration of the puck

a= u g = 0.15*9.8 = 1.47 m/s^2

using 1st equation of motion

v = u - at

18.865 = 20 - 1.47 t

t = 0.772 seconds

========

do comment in case any doubt, will reply for sure.. Goodluck